I have to find the area and perimeter of a rectangle. On the top of the rectangle is d+10 and on the right side of the rectangle just has a single d. How do I find these without having the length and width numbers?
Answer by ankor@dixie-net.com(22740)

I have to find the area and perimeter of a rectangle.
On the top of the rectangle is d+10 and on the right side of the rectangle just has a single d.
:
:
d has to be part of the answer
:
The perimeter:
p = 2(d+10) + 2d
p = 2d + 20 + 2d
p = 4d + 20
:
The area
A = d*(d+10)
A = d^2 + 10d